If a+b=2 and b*a= −1, then a^2 + b^2 =
(A)4
(B)5
(C)6
(D)8
(E)10

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

If a+b=2 and b*a= −1, then a^2 + b^2 =
(A)4
(B)5
(C)6
(D)8
(E)10
================
With only 5 to choose from, try each one.

Answer by ikleyn(52780)

.

Since a + b = 2, square both sides to get


    (a + b)^2 = 4

    a^2 + 2ab + b^2 = 4


Next, replace ab by -1, since it is given.  You will get


    a^2 -2 + b^2 = 4,


which implies


    a^2 + b^2 = 4 + 2 = 6.


ANSWER.  Under the given conditions,  a^2 + b^2 = 6.

Solved.


/\/\/\/\/\/\/\/\/\/\/\/


-------------
Comment from student:  Thanks a lot Ikleyn.  What is the link of your website?  I can share.
-------------


My response :   This forum is constructed in the way that i  CAN  NOT  identify from whom I got the message/(the  "comment").

Therefore,  when I get the comment from the visitor in this way,  I simply  DON'T  KNOW  to whom to answer.

If you really want to get the response from me to your message/"comment",
then place the problem's ID number to your message/comment.

It is the only way to support a  two-sides/(two-ways)  dialog.